one exterior angle of a regular polygon measures 30 degrees. what is the sum of the polygons interior angles measures?
Answer by stanbon(75887)

One exterior angle of a regular polygon measures 30 degrees. what is the sum of the polygons interior angles measures?
----
# of exterior angles = 360/30 = 12
---
Each interior angle = 180-30 = 150
-----
Total interior angle sum = 12*150 = 1800 degrees
----------------------
Cheers,
Stan H.
===========